Imagine a right circular cone with altitude H and slant height R. Let the radius of

the circular base be r. Imagine that the apex angle of the cone is "\theta" . Now slit the

cone open by cutting along the slant height from the base to the apex. Unwrap the

slit cone to reveal a circular sector. Find the central angle of this sector, "\alpha" , as a

function of "\theta" .


1
Expert's answer
2021-11-02T18:04:00-0400

the radius of circular sector = slant height R

surface area of cone = area of a sector


surface area of cone:

"A=πr(r+\sqrt{H^2+r^2})=\pi r(r+R)"


area of a sector:

"(\alpha/360º) \cdot πR^2=\pi r(r+R)"


for apex angle of the cone:

"sin(\theta/2)=r/R"


then:

"(\alpha/360º) = sin(\theta/2)(sin(\theta/2)+1)"


"\alpha= sin(\theta/2)(sin(\theta/2)+1)\cdot 360\degree"